# Hallucination, reliability, and the role of generative AI in science

### v1

@@ Abstract
Generative AI is increasingly used in scientific domains, from protein folding to climate modeling. But these models produce distinctive errors known as “hallucinations”—outputs that are incorrect yet superficially plausible. Worse, some arguments suggest that hallucinations are an inevitable consequence of the mechanisms underlying generative inference. Fortunately, such arguments rely on a conception of hallucination defined solely with respect to internal properties of the model, rather than in reference to the empirical target system. This conception fails to distinguish epistemically benign errors from those that threaten scientific inference. I introduce the concept of corrosive hallucination to capture the epistemically troubling subclass: misrepresentations that are substantively misleading and resistant to systematic anticipation. I argue that although corrosive hallucinations do pose a threat to scientific reliability, they are not inevitable. Scientific workflows such as those surrounding AlphaFold and GenCast, both of which serve as case studies, can neutralize their effects by imposing theoretical constraints during training, and by strategically screening for errors at inference time. When embedded in such workflows, generative AI can reliably contribute to scientific knowledge.
footnotetext: This is a preprint of a draft paper intended for journal submission. Comments are very welcome.
@@
Word count: 8,070
Keywords: generative AI, hallucination in AI, corrosive hallucination, scientific AI, AI reliability, reliabilist epistemology, AI epistemology
@@ 1 Hallucination as a threat to reliability
There is considerable confusion about what hallucinations are. The term itself invites a misleading comparison to human perceptual hallucinations. Generative AI models do not consciously perceive the world, let alone misperceive it. Nevertheless, the real problem lies not with the word, but with the errors themselves. Since the term is already widely used, and since insisting on a replacement would only introduce a cognitively costly neologism into an already difficult discussion, I will just stick with it.
@@
As a first pass, hallucinations can be characterized as errors that are not merely inherited from the training data, but are, in some sense, produced by the model itself. This claim is vague, but also revealing. After reading it, your first thought might well be that any model prone to hallucination ought not to be trusted. And indeed, the epistemic risks introduced by hallucinations are serious. AlphaFold 3, among the most celebrated generative AI models in the natural sciences, has been shown to generate detailed molecular structures where none exist [Abramson et al., 2024]. Generative adversarial networks (GANs), widely used in medical imaging, have been found to introduce phantom anomalies—a fracture-like line appearing in an unbroken bone, or a lesion in what is, in fact, perfectly healthy tissue [Shin et al., 2021]. These are not just rounding errors. Undetected, hallucinations can lead researchers and clinicians toward serious mistakes in inference and decision-making.
@@
In fact, the epistemic challenges posed by hallucinations run deeper than these examples suggest. There is reason to think that hallucinations are inevitable byproducts of the mechanisms of generative inference. The intuition behind this claim is that training such models involves a fundamental tradeoff between novelty and reliability [Sajjadi et al., 2018; Sinha et al., 2023; Xu et al., 2024]. A model constrained to strictly mirror its training data may be reliable but incapable of generating novel insights. Allowing a model to extrapolate, by contrast, enables novelty but invites fabrication.
@@
Another reason that hallucinations threaten reliability is that they are sometimes difficult to detect. [Ji et al., 2023; Bubeck et al., 2023]. This is not always the case. When we have thorough background knowledge of the target phenomenon, hallucinations can be easy to spot. For example, earlier versions of DALL-E and Stable Diffusion often generated images of human hands with six fingers [Wang et al., 2024]. But scientific AI operates at the frontiers of human knowledge, where error detection is intrinsically more difficult. Where our epistemic resources are weakest, errors are most likely to go undetected. And the longer they remain undetected, the more they threaten to derail any decision-making processes based on model outputs.
Drawing these observations together, it seems that hallucinations are, in at least some cases, substantive, inevitable, and difficult to detect. It is as if hallucinations were tailor-made—perhaps by a Cartesian demon—to undermine the reliability of scientific inference.
What can be done to mitigate the problem? That question is difficult to answer, in part because deep learning models are epistemically opaque [Humphreys, 2009; Creel, 2020]. In traditional closed-form models, evidence of reliability is often grounded in knowledge of how parameters relate to the properties of the target system. In DNNs, by contrast, it is unclear whether individual parameters represent anything at all. This lack of interpretability makes it difficult to assess whether a model tracks the underlying structure of the world. As Sullivan [2022] emphasizes, genuine scientific understanding requires more than predictive accuracy: it requires that the model’s internal representations align in some meaningful way with the causal structure of the target phenomenon. If we are to justify the use of opaque, DNN-based models, we must find alternative grounds for assessing their epistemic reliability.
If we do not fully understand the mechanisms behind generative AI, how might we justify its use? Reliabilist epistemology [Goldman, 1979; Lyons, 2019] offers a straightforward answer: we observe its track record. Instead of explaining why a model succeeds, we infer its reliability from past performance. This approach, which Duede [2023] calls brute inductivism, reduces scientific epistemology to an accounting exercise. Suppose a model achieves high accuracy on benchmarks or aligns well with historical data. Researchers then infer—perhaps naively—that the model will be reliable in future applications. But as Duede’s unflattering label suggests, brute inductivism is an inherently precarious strategy. Past success offers no guarantee of future performance, particularly in novel settings [Grote et al., 2024]. If trust in generative AI rests on nothing more than inductive bookkeeping, it risks collapsing as soon as the model is pushed beyond its training distribution.
Despite these mutually reinforcing difficulties of hallucination and epistemic opacity, rejecting generative AI outright is not a serious option. These models have already demonstrated their ability to outperform traditional approaches in all sorts of important predictive tasks. The real challenge, then, is not whether generative AI should be used in science, but how it can be used responsibly.
Addressing this challenge requires greater clarity about what counts as a hallucination. The concept is often defined formally, and while formal definitions are always welcome, they sometimes obscure an important empirical distinction. In the case of hallucination, such definitions lump together epistemically threatening errors with those that are largely benign. To address the problem, I introduce a new concept, corrosive hallucination, which is specifically designed to isolate those hallucinations that pose a genuine threat to the reliability of science. It will turn out, however, that when we focus specifically on corrosive hallucinations, arguments for the inevitability of hallucination collapse. Mitigation is within our grasp after all.
@@
To illustrate this, I examine two case studies: AlphaFold 3, which predicts molecular structures, and GenCast, which generates probabilistic weather forecasts. These models operate in entirely different scientific domains—one at the scale of molecules, the other at the scale of planetary weather systems. Nevertheless, both mitigate hallucinations by embedding theoretical constraints and uncertainty management strategies directly into their modeling architectures. These principles do not eliminate model-generated error entirely, but they show how, despite the distinctive challenges posed by generative AI, such errors can be effectively managed.
Crucially, these design principles are neither automatic nor inevitable. They emerge from carefully managed scientific workflows, and their effectiveness depends on deliberate design and maintenance. By articulating the rationale behind these strategies, I aim to clarify how generative AI can be integrated into scientific practice without unduly compromising reliability.
2 Generative AI and the inevitability of hallucination
@@ 2.1 What is generative AI?
Two contrasts help clarify what makes generative AI distinctive. First, unlike classification models, which learn a conditional distribution $P(Y\mid X)$ over discrete labels $Y$, generative models aim to learn the full distribution $P(X)$, enabling them to produce novel samples that extend the distribution in coherent ways [Kingma and Welling, 2013; Goodfellow et al., 2014; Buckner, 2024]. Second, unlike classical generative statistical models such as Poisson processes or Markov chains, which generate data from predefined parametric distributions [Grimmett and Stirzaker, 1992], generative AI models learn latent representations that capture complex, often idiosyncratic statistical structure [Rezende et al., 2014; Yang et al., 2023]. This capacity makes them uniquely valuable for domains where explicit theory remains incomplete, such as materials science or drug discovery.
2.2 On the inevitability of hallucination
@@
Unlike autoregressive architectures, which are well suited to sequential data such as text, many scientific generative models are designed to preserve global coherence across high-dimensional structures. Diffusion models, in particular, generate outputs through a process of iterative refinement, progressively denoising a sample over multiple steps [Song et al., 2021]. Variational autoencoders (VAEs) and generative adversarial networks (GANs), though architecturally distinct, pursue the same end: to reconstruct complex global structure from sparse data. These models are typically applied in domains where long-range dependencies span multiple spatial or structural dimensions—protein folding, weather dynamics, material synthesis. I focus on these architectures for two reasons. First, they have figured centrally in some of the most celebrated successes of generative AI in the natural sciences. Second, because their outputs are not merely large but genuinely high-dimensional—structured across space, geometry, or topology—the detection of hallucination poses distinct challenges. Unlike language models, which produce long sequences of discrete tokens, these models (often) generate high-dimensional outputs, in which hallucination detection is intrinsically more difficult. In what follows, therefore, I borrow some ideas from inevitability arguments developed for language models to this broader class of scientific models.
One kind of argument is broadly information-theoretic. The idea is that generative AI models do not contain enough information to represent complex empirical distributions accurately. To see this, consider the size of the output space relative to the model’s internal parameter space. Generative models operate in high-dimensional output spaces, with far more possible configurations than any dataset can sample faithfully. For example, a 12-megapixel image with 256 intensity levels per channel has $10^{86,000,000}$ possible configurations. A 100-amino acid protein has $10^{130}$ possible sequences, not counting conformational variants [Dryden et al., 2008]. Meteorological states evolve across millions of interacting variables. Even the largest training sets cover only a vanishing fraction of these spaces. Moreover, models compress these sparse samples into relatively small parameter sets. A protein diffusion model may train on a few hundred thousand examples, but must generalize across $10^{100}$ possible sequences and conformations. This compression all but ensures that many outputs will be generated in regions where the training data provides little guidance. And where the training data provides little guidance, hallucination is inevitable.
@@
Generative models tend to capture short-range dependencies more faithfully than long-range ones. This reflects a basic statistical fact: the nearer the elements, the clearer the pattern. Local structures—bond angles in molecules, temperature gradients in weather fields—recur with high signal and low variation. Long-range dependencies, by contrast, are more easily obscured by noise. As a result, generative models, whether built on diffusion processes or transformers, often produce outputs that are locally plausible but globally flawed. A protein may contain chemically sound fragments yet fold into an unstable conformation. A weather forecast may model regional dynamics with precision while violating large-scale conservation laws. Large language models exhibit a parallel tendency: they produce coherent sentences and paragraphs that fail to cohere at the level of extended argument. In each case, local plausibility masks more distributed structural flaws.
@@ 2.3 Hallucination in diffusion models, and a proposed solution
Aithal et al. train diffusion models on synthetic datasets specifically designed to make the structure of the data manifold transparent. In one experiment, the training data is sampled from a mixture of eight well-separated Gaussians, with each point drawn from a single mode. In another, they use binary 10$\times$10 grids, constrained so that exactly half the cells are activated according to simple structural rules. The purpose of these setups is to ensure that the generative principles underlying the training data are fully known. This allows them to test whether a diffusion model can learn those principles without producing spurious outputs.
@@
This filtering solution is entirely reasonable when applied to synthetic datasets designed to test model behavior, rather than to represent an external phenomenon of scientific interest. However, in a scientific setting, our primary goal is to acquire information about the nature of the target. Specifically, the aim is to discover hidden structure that is not explicitly represented in the training data. But here, we encounter a fundamental difficulty with the conception of hallucination we have employed thus far: a generated output that falls between known modes may signal either a modeling error or a discovery—an instance of structure that the training data failed to make explicit. In the context of scientific inquiry, then, Aithal et al. [2025]’s solution is too conservative: it eliminates precisely those cases where the most interesting knowledge might emerge. If hallucination were simply a matter of deviation from training data, then every output that deserves to be called a genuine discovery would, ipso facto, count as a hallucination, and nearly all of those would get filtered out.
@@
3 An epistemic account of hallucination
3.1 Data and phenomena
I want to zoom out briefly and draw a parallel between the data-centric attitude that seems prevalent in AI today and a similar attitude that prevailed in 20th-century philosophy of science. The logical empiricists (especially Carnap [1928] in the Aufbau) viewed science as the reconstruction of observational data. That view withered under criticism, but the underlying idea that theories earn legitimacy only by recovering or predicting patterns in data of some sort seems to have been widely accepted well past the middle part of the twentieth century. But as Bogen and Woodward (1988) forcefully argued, this outlook fails to account for the constructed nature of data. Because most of the data sets that scientists work with are shaped by the contingencies of measurement techniques and experimental design, scientific reasoning necessarily involves questions about how data can sometimes mislead us about the nature of the target phenomenon. To put this thought in slogan form, the data are a means to an end, rather than an end in themselves.
@@ 3.1 Data and phenomena
3.2 Hallucination and the anticipation of error
With this picture in place, we are now in position to appreciate the epistemic threat that hallucination poses. Because existing discussions rarely address the epistemic dimension of hallucination, there is no established term that picks out all and only threatening errors. The task, then, is not to analyze an existing concept but to construct one. The goal is to work out what properties a hallucination-like error would have to have, such that, if left unchecked, it would undermine the reliability of scientific inference. I will call errors of this kind corrosive hallucinations.
To develop this concept, we must first ask how hallucinations relate to the empirical target phenomenon. [^3] My emphasis on the empirical character of the target system has the consequence that the present analysis does not apply to models in pure mathematics. This is not meant to imply that similar issues cannot crop up in that setting. But I do think that the application of deep learning to pure mathematics raises distinct philosophical questions that deserve their own analysis. With Figure 1 in mind, we can say that a hallucination is an output that misrepresents the target. But the question of whether some data structure counts as a misrepresentation is subtle. What counts as a misrepresentation will depend on the interpretive conventions of the relevant discipline, and on the interpretive habits of the relevant scientists.
The website thesecatsdonotexist.com provides a whimsical but intuitively compelling illustration of this point. The website produces realistic images of cats using a StyleGAN model (a type of generative adversarial network trained to synthesize highly realistic images). If we interpret the outputs as images of particular, spatiotemporally located cats, then the images are hallucinations. As the url suggests, these cats do not exist. If, however, we interpret them as images that merely resemble cats, then they are not hallucinations. They provide us with accurate knowledge of the statistical properties of cat appearances.
Scientific modeling is, in a sense, similar. Researchers use model outputs to draw inferences about the target phenomenon. It is this act of interpretation—treating the output as evidence for a claim about the target—that determines whether a misrepresentation occurs. Some hallucination-like errors are harmless or even illuminating, depending on how the output is interpreted. In fact, this point applies more forcefully in scientific cases than in the case of cat images, because only the latter is tightly constrained by the operation of the human perceptual system.
@@ 3.2 Hallucination and the anticipation of error
I said that hallucinations are misrepresentations. But misrepresentation is not necessarily a bad thing. Many scientific misrepresentations are introduced deliberately, as idealizations. Philosophers have long noted that models include systematic distortions—treating gases as point particles or assuming frictionless planes—to render phenomena tractable [Weisberg, 2007; Strevens, 2016]. These are strategic distortions designed to facilitate inference. Hallucinations, by contrast, are unintentional artifacts of generative processes.
But even this narrower class of unintentional misrepresentation does not yet suffice as an analysis of hallucination. Scientific models frequently produce errors, whether due to approximation, numerical instability, or incomplete training data, that are unintentional but still manageable. Not every unintentional error threatens the reliability of inference. To distinguish hallucination from other forms of benign misrepresentation, we need an additional criterion. That criterion is resistance to anticipation.
Here I use the term “anticipation” broadly. It refers to any strategy that enables us to foresee the conditions under which errors are more likely. If we can do that, we gain some control over a model’s reliability. We can adjust the model itself to prevent further errors of the same type, reinterpret the outputs to account for known failure modes, or narrow the model’s domain of application to avoid settings in which such errors are likely.
This kind of anticipation is far more difficult in generative models than in classification models, where the task is often tractable. In classification, outputs are discrete and error types are well-defined. This makes it possible to systematically probe model performance across controlled variations in input space. Researchers can vary input features, observe changes in output, and generalize about which kinds of inputs lead to failure. For example, if a radiology classifier performs poorly on underexposed images, this fact can be discovered by manipulating illumination conditions and observing the effect on predictions. The underlying causal structure of the model’s behavior—how inputs contribute to specific types of errors—is accessible to empirical investigation. Attribution methods like SHAP [Lundberg and Lee, 2017] or saliency maps [Simonyan et al., 2014] leverage this structure to connect input features to model decisions.
With generative models, this kind of systematic circumscription of the input space is radically more difficult. Because outputs are high-dimensional—images, molecular geometries, weather fields, and so on—there are countless ways for them to be wrong. Moreover, failure modes often interact: fixing one flaw may introduce another. This complexity makes it hard to trace any single failure back to a specific cause in the input. The stable, localizable causal dependencies that support generalization in simpler models begin to break down.
Now that we have described the epistemic threat posed by hallucination, we can use it to construct a definition.
An error counts as a corrosive hallucination iff it is both epistemically disruptive and resistant to anticipation.
@@
    Epistemically disruptive. An error is epistemically disruptive if it is both substantive and non-strategic.
    1. (a)
      Substantive: It alters downstream reasoning—it changes conclusions, decisions, or inferential trajectories. Crucially, this significance depends on interpretation: an output is only epistemically disruptive when it is taken as evidence for a particular claim about the target phenomenon.
    2. (b)
      Non-strategic: It does not arise from an intentional simplification designed to aid understanding. Idealizations—such as modeling gases as point particles—introduce systematic distortions but serve an epistemic function. Corrosive hallucinations, by contrast, distort without enlightening.
@@
    Resistant to anticipation. An error exhibits resistance to anticipation when there is no known method for systematically identifying the conditions under which it is more or less likely to occur. In particular, researchers lack tools for mapping its frequency across regions of input space or tracing it to identifiable features of the model’s behavior. As a result, such errors resist localization, prediction, and control.
This definition names the threat and marks the first step toward addressing it. Happily, not all errors produced by generative models are corrosive. Some deviations from ground truth are insignificant; others, though epistemically relevant, are anticipated and therefore safely bracketed. The danger arises only when hallucinations are both epistemically disruptive and systematically resistant to anticipation. The remainder of this paper considers how even these most troubling errors can be managed effectively.

### v2

# From Hallucination to Reliability: Generative Modeling and the Structure of Scientific Inference

@@ Abstract
Generative AI increasingly supports scientific inference, from protein structure prediction to weather forecasting. Yet its distinctive failure mode, hallucination, raises epistemic alarm bells. I argue that this failure mode can be addressed by shifting from data-centric to phenomenon-centric assessment. Through case studies of AlphaFold and GenCast, I show how scientific workflows discipline generative models through theory-guided training and confidence-based error screening. These strategies convert hallucination from an unmanageable epistemic threat into bounded risk. When embedded in such workflows, generative models support reliable inference despite opacity, provided they operate in theoretically mature domains.
@@
Word count: 8,331 Keywords: generative AI, hallucination in AI, scientific AI, AI reliability, reliabilist epistemology, AI epistemology, protein structure prediction, weather forecasting
@@ 1 Hallucination as a threat to reliability
As a first pass, hallucinations [^2] The term invites a misleading comparison to human perceptual experience. Generative AI models do not consciously perceive the world, let alone misperceive it. Nevertheless, since the term is already widely used in the technical literature, insisting on a replacement would only introduce a cognitively costly neologism into an already difficult discussion. can be characterized as errors that are not merely inherited from the training data, but are, in some sense, produced by the model itself. This claim is substantive: not all errors count as hallucinations. A corrupted measurement or a mislabeled datapoint is an error, but not a hallucination. I return to a more precise analysis in Section 3, but even this initial sketch helps explain why the phenomenon demands attention. It is natural to worry that any model prone to hallucination may not be trustworthy—and indeed, the epistemic risks are serious. AlphaFold 3, among the most celebrated generative models in science, has been shown to produce detailed molecular structures where none exist [Abramson et al., 2024]. GANs used in medical imaging have introduced phantom anomalies—a fracture-like line in an unbroken bone, or a lesion in healthy tissue [Shin et al., 2021].
@@
These are not just rounding errors. Undetected, hallucinations can lead researchers and clinicians toward serious mistakes in inference and decision-making. In fact, the epistemic challenges posed by hallucinations run deeper than these examples suggest. There is reason to think that hallucinations are inevitable byproducts of the mechanisms of generative inference. The intuition behind this claim is that training such models involves a fundamental tradeoff between novelty and reliability [Sajjadi et al., 2018; Sinha et al., 2023; Xu et al., 2024]. A model constrained to strictly mirror its training data may be reliable but incapable of generating novel insights. Allowing a model to extrapolate, by contrast, enables novelty but invites fabrication.
@@
Another reason that hallucinations threaten reliability is that they are sometimes difficult to detect [Ji et al., 2023; Bubeck et al., 2023]. This is not always the case. When we have thorough background knowledge of the target phenomenon, hallucinations can be easy to spot. For example, earlier versions of DALL-E and Stable Diffusion often generated images of human hands with six fingers [Wang et al., 2024]. But scientific AI operates at the frontiers of human knowledge, where error detection is intrinsically more difficult. Where our background knowledge is weakest, errors are most likely to go undetected. And the longer they remain undetected, the more they threaten to derail any decision-making processes based on model outputs.
Drawing these observations together, it seems that hallucinations are, in at least some cases, substantive, inevitable, and difficult to detect. Worse still, deep learning models are epistemically opaque [Humphreys, 2009; Creel, 2020]. In traditional closed-form models, evidence of reliability is often grounded in knowledge of how parameters relate to the properties of the target system. When errors arise, they can be traced to specific parameters and corrected. In DNNs, by contrast, it is unclear whether individual parameters represent anything at all.
This combination of opacity and model-generated error creates what I will call the diagnostic problem: once we discover a hallucination downstream, how can we use that knowledge to systematically improve the model? In normal scientific practice, error correction is iterative. When a model produces questionable results, researchers trace the error back to specific parameters, adjust them based on their representational role, and thereby reduce the probability of similar errors in future applications. Opacity blocks precisely this diagnostic workflow. Without knowing which parameters are responsible for an error, we cannot learn from our mistakes in the systematic way science typically demands. If we cannot check the representational fidelity of individual parameters, how might we justify using these models at all? Reliabilist epistemology [Goldman, 1979; Lyons, 2019] offers a straightforward answer: we observe its track record. Instead of explaining why a model succeeds, we infer its reliability from past performance. This approach, which Duede [2023] calls brute inductivism, reduces scientific epistemology to an accounting exercise.
Suppose a model achieves high accuracy on benchmarks or aligns well with historical data. Researchers then infer—perhaps naively—that the model will be reliable in future applications. But as Duede’s unflattering label suggests, brute inductivism is an inherently precarious strategy. Past success offers no guarantee of future performance, particularly in novel settings [Grote et al., 2024]. Nevertheless, these models have already demonstrated their ability to outperform traditional approaches in all sorts of important predictive tasks. The real challenge, then, is not whether generative AI should be used in science, but how it can be used responsibly.
Addressing this challenge requires greater clarity about what counts as a hallucination. Existing definitions, whether formal or informal, tend to evaluate model outputs primarily by their relationship to training data. But, as I will argue in Section 3, this data-centric approach forecloses the very solutions that make successful applications possible. If we assess hallucinations by their deviation from training data, we are led toward filtering strategies that would eliminate precisely those outputs where genuine scientific discovery occurs. What matters for scientific reliability is not whether outputs deviate from training data, but whether they misrepresent the target phenomena we aim to understand. Shifting to this phenomenon-centric view reveals that many outputs flagged as hallucinations under data-centric definitions pose no genuine epistemic threat, while others—those that cannot be reliably detected or filtered—demand more careful management.
@@
To illustrate how scientists address the diagnostic problem, I examine two case studies: AlphaFold 3, which predicts molecular structures, and GenCast, which generates probabilistic weather forecasts. These models operate in entirely different scientific domains—one at the scale of molecules, the other at the scale of planetary weather systems. Nevertheless, both mitigate hallucinations by embedding theoretical constraints and uncertainty management strategies directly into their modeling architectures. These principles do not eliminate model-generated error entirely, but they show how, despite the distinctive challenges posed by generative AI, such errors can be effectively managed. Crucially, these design principles are neither automatic nor inevitable. They emerge from carefully managed scientific workflows, and their effectiveness depends on deliberate design and maintenance. By articulating the rationale behind these strategies, I aim to clarify how generative AI can be integrated into scientific practice without unduly compromising reliability.
2 On the inevitability of hallucination in generative AI
@@ 2.1 What is generative AI?
Two contrasts help clarify what makes generative AI distinctive. First, unlike classification models, which learn to represent the conditional distribution $P(Y\mid X)$ over discrete labels $Y$, generative models aim to learn a representation of the full distribution $P(X)$, enabling them to produce novel samples that extend the distribution in coherent ways [Kingma and Welling, 2013; Goodfellow et al., 2014; Buckner, 2024]. Second, unlike classical generative statistical models such as Poisson processes or Markov chains, which generate data from predefined parametric distributions [Grimmet and Sterzaker, 1992], generative AI models learn latent representations that capture complex, often idiosyncratic statistical structure [Rezende et al., 2014; Yang et al., 2023]. This capacity makes them uniquely valuable for domains where explicit theory remains incomplete, such as materials science or drug discovery.
2.2 Inevitability arguments
@@
Unlike autoregressive architectures, which are well suited to sequential data such as text, many scientific generative models are designed to preserve global coherence across high-dimensional structures. Diffusion models, in particular, generate outputs through a process of global iterative refinement, progressively denoising a sample over multiple steps [Song et al., 2021]. Variational autoencoders (VAEs) and generative adversarial networks (GANs), though architecturally distinct, pursue the same end: to reconstruct complex global structure from sparse data. These models are typically applied in domains where long-range dependencies span multiple spatial or structural dimensions—protein folding, weather dynamics, material synthesis. I focus on these architectures for two reasons. First, they have figured centrally in some of the most celebrated successes of generative AI in the natural sciences. Second, because their outputs are not merely large but genuinely high-dimensional—structured across space, geometry, or topology—the detection of hallucination poses distinct challenges. Unlike language models, which produce long sequences of discrete tokens, these models (often) generate high-dimensional outputs, in which hallucination detection is intrinsically more difficult. In what follows, therefore, I borrow some ideas from inevitability arguments developed for language models to this broader class of scientific models.
One kind of argument is broadly information-theoretic. The idea is that generative AI models do not contain enough information to represent complex empirical distributions accurately. To see this, consider the size of the output space relative to the model’s internal parameter space. Generative models operate in high-dimensional output spaces, with far more possible configurations than any dataset can sample faithfully. For example, a 12-megapixel image with 256 intensity levels per channel has $10^{86,000,000}$ possible configurations. A 100-amino acid protein has $10^{130}$ possible sequences, not counting conformational variants [Dryden et al., 2008]. Meteorological models, for example, must track millions of degrees of freedom—combinations of pressure, temperature, humidity, and other variables across thousands of spatial and temporal points. Even the largest training sets cover only a vanishing fraction of these spaces. Moreover, models compress these sparse samples into relatively small parameter sets. A protein diffusion model may train on a few hundred thousand examples, but must generalize across $10^{100}$ possible sequences and conformations. This compression all but ensures that many outputs will be generated in regions where the training data provides little guidance. And where the training data provides little guidance, hallucination is inevitable.
@@
Generative models tend to capture short-range dependencies more faithfully than long-range ones. This reflects a basic statistical fact: the nearer the elements, the clearer the pattern. Local structures such as bond angles in molecules or temperature gradients in weather fields recur with high signal and low variation. Long-range dependencies, by contrast, are more easily obscured by noise. They often involve more subtle or indirect interactions, and models may lack both the capacity to represent them and the training data to learn them reliably. As a result, generative models, whether built on diffusion processes or transformers, often produce outputs that are locally plausible but globally flawed. A protein may contain chemically sound fragments yet fold into an unstable conformation. A weather forecast may model regional dynamics with precision while violating large-scale conservation laws. Large language models exhibit a parallel tendency: they produce coherent sentences and paragraphs that fail to cohere at the level of extended argument. In each case, local plausibility masks more distributed structural flaws.
@@ 2.3 Hallucination in diffusion models, and a proposed solution
Aithal et al. train diffusion models on synthetic datasets specifically designed to make the structure of the data manifold transparent. In one experiment, the training data is sampled from a mixture of eight well-separated Gaussians, with each point drawn from a single mode. In another, they use binary $\times10$ grids, constrained so that exactly half the cells are activated according to simple structural rules. The purpose of these setups is to ensure that the generative principles underlying the training data are fully known. This allows them to test whether a diffusion model can learn those principles without producing spurious outputs.
@@
This filtering solution is entirely reasonable when applied to synthetic datasets where the data-generating process is both simple and fully specified, and designed to test model behavior rather than to represent an external phenomenon of scientific interest. However, in a scientific setting, our primary goal is to acquire information about the nature of the target. Specifically, the aim is to discover hidden structure that is not explicitly represented in the training data. But here, we encounter a fundamental difficulty with the conception of hallucination we have employed thus far: a generated output that falls between known modes may signal either a modeling error or a discovery—an instance of structure that the training data failed to make explicit. In the context of scientific inquiry, then, Aithal et al. [2025]’s solution is too conservative: it eliminates precisely those cases where the most interesting knowledge might emerge. If hallucination were simply a matter of deviation from training data, then every output that deserves to be called a genuine discovery would, ipso facto, count as a hallucination, and nearly all of those would get filtered out.
@@
3 Rethinking hallucination
Before developing an improved analysis of hallucination, I want to zoom out briefly and draw a parallel between the data-centric attitude that seems prevalent in AI today and a similar attitude that prevailed in 20th-century philosophy of science. The logical empiricists (especially Carnap [1928] in the Aufbau) viewed science as the reconstruction of observational data. That view withered under criticism, but the underlying idea that theories earn legitimacy only by recovering or predicting patterns in data of some sort seems to have been widely accepted well past the middle part of the twentieth century. But as But as Bogen and Woodward [1988] forcefully argued, this outlook fails to account for the constructed nature of data. Because most of the data sets that scientists work with are shaped by the contingencies of measurement techniques and experimental design, scientific reasoning necessarily involves questions about how data can sometimes mislead us about the nature of the target phenomenon. To put this thought in slogan form, the data are a means to an end, rather than an end in themselves.
@@ 3 Rethinking hallucination
In the introduction I said that a hallucination is an error that is produced by the model (rather than one that is inherited from the training data.) An error is a deviation from some standard, and the picture above makes it clear that the target phenomenon, rather than the training data, is the relevant standard. But now we should also ask: what kind of deviations count as errors? That question has no straightforward answer because what counts as an error depends in part on the interpretive practices of the relevant scientific community.
To see why interpretation matters, consider a non-scientific example. The website thesecatsdonotexist.com produces realistic images of cats using a StyleGAN model trained on photographs of real cats. Now suppose you were learning about cats from a sequence of images that included both real photographs and outputs from this model. According to the canonical interpretive scheme for photographs, according to which they depict particular, spatiotemporally located individuals, the StyleGAN images count as misrepresentations. However, if you can reliably identify which images come from the StyleGAN, these misrepresentations need not stand in the way of the acquisition of new knowledge. The synthetic images still provide accurate information about the statistical properties of cat-like appearances, even though they fail as photographic representations of particular animals.
The crucial point is that an output can be a misrepresentation according to the standard interpretive scheme for its domain, and nevertheless remain epistemically benign as long as users can detect it and adjust their interpretive stance accordingly. When mixed with genuinely photographic images, the StyleGAN images are hallucinations, but, even if they differ systematically from genuine photos in ways that are hard to see, they need not cause any false beliefs about cats. [^4] This is why we can coherently talk about detecting, flagging, and filtering hallucinations without contradiction. Hallucination is defined by misrepresentation according to canonical interpretive schemes, not by capacity to produce false belief. Successful detection simply prevents epistemically harmful misrepresentations from propagating through the workflow. In scientific contexts, similarly, an output might misrepresent the target phenomenon according to standard domain conventions, and yet remain benign or even useful if researchers can identify it and interpret it appropriately.
@@
This puts us in position to distinguish hallucinations from another familiar category of scientific misrepresentation: idealizations. Scientific models include systematic distortions—treating gases as point particles or assuming frictionless planes—to render phenomena tractable [Weisberg, 2007; Strevens, 2016]. These are strategic distortions, deliberately introduced to facilitate inference. Idealizations involve misrepresentation, but they operate within a well-understood interpretive framework where the distortions are controlled and their effects on downstream inference are anticipated. Hallucinations, by contrast, are non-strategic misrepresentations—unintentional artifacts of the generative process that must be either filtered out or managed through interpretive reframing.
With these distinctions in place, we can now define hallucinations for scientific applications of generative AI. A hallucination is a generative AI model output that satisfies three conditions:
@@
    It counts as a misrepresentation of the target system according to the canonical interpretive scheme for outputs of that kind.
@@
    It is non-strategic: an unintended artifact of the generative process rather than a deliberate idealization.
  3. 3.
    Its misrepresentational status was produced by the model’s generative activity, rather than having been inherited from the training data.
 [^5] Because this definition is restricted to generative AI models, it implicitly incorporates the high-dimensional, structured nature of their outputs. As discussed in Section 2.1, this complexity is central to understanding both why hallucinations are inevitable and why they pose distinctive epistemic challenges in scientific contexts.
This definition is intended to pick out the class of hallucinations more accurately than existing alternatives, but it is also structured to help analyze how scientists address the diagnostic problem introduced earlier. The three-part definition directs our attention to the right questions: (1) What interpretive scheme governs scientific practice in this domain? (2) How do workflows distinguish strategic from non-strategic misrepresentations? (3) What mechanisms prevent model-generated errors from propagating? Our case studies show that when these questions are answered carefully, the diagnostic problem becomes tractable. Recall that the problem arises from the combination of opacity and model-generated error: once we discover a hallucination downstream, how can we use that knowledge to improve reliability? The traditional approach—localizing the guilty parameter and adjusting it—is unavailable. But as we will see, scientists work around this limitation not by making models transparent, but by embedding them in workflows that manage error at the level of outputs rather than parameters.